OpenStudy (anonymous):

i give metals and i will be your fan. What is the value of the expression 7^3?

OpenStudy (anonymous):

@SolomonZelman @Nnesha

Nnesha (nnesha):

the exponents tells how many times you should multiply the base 2^4 = 2 times 2 times 2 times 2

OpenStudy (adilalvi):

343

OpenStudy (anonymous):

the 4 is a negative

OpenStudy (anonymous):

so -343

Nnesha (nnesha):

ohh then you should apply the exponent rule \[\huge\rm 7^{-3}\] like this ?

OpenStudy (anonymous):

ya

Nnesha (nnesha):

\[\huge\rm x^{-m}=\frac{ 1 }{ x^m }\] flip the fraction when u flip it , the sign of the exponent would change

OpenStudy (adilalvi):

yaah if its negitive then it will be -343

OpenStudy (anonymous):

so \[1/343\]

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

can you help me with more

Nnesha (nnesha):

i'll try :=)

OpenStudy (anonymous):

What is the value of the expression ? 1/5^-5

OpenStudy (anonymous):



Nnesha (nnesha):

okay apply the same rule!

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \[\huge\rm x^{-m}=\frac{ 1 }{ x^m }\] flip the fraction when u flip it , the sign of the exponent would change \(\color{blue}{\text{End of Quote}}\) exponent is negative so ^^^^^^

OpenStudy (anonymous):

3125 is the anwser

Nnesha (nnesha):

well don't use the calculator.

OpenStudy (anonymous):

1/3125

Nnesha (nnesha):

no flip the fraction

OpenStudy (anonymous):

how can you show me

Nnesha (nnesha):

okay if i have \[\huge\rm 2^{-2} =\frac{ 1 }{ 2^2 }\] i would flip the fraction and change the sign of exponent

OpenStudy (anonymous):

so it would just be 3125

Nnesha (nnesha):

yes right i guess you don't have to find 5^5

Nnesha (nnesha):

are there answer choices ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

okay then that's right

OpenStudy (anonymous):

A. 1/3125 B.1/25 C.25 D.3125

Nnesha (nnesha):

ookay. then ur answer is correct :=)

OpenStudy (anonymous):

so D

Nnesha (nnesha):

yes

OpenStudy (anonymous):

What is the value of the expression

Nnesha (nnesha):

rewrite 4 in terms of base 2

OpenStudy (anonymous):

here are the options

OpenStudy (anonymous):

so 4/3

Nnesha (nnesha):

no.. here is an example rewrite 9 in terms of base 3 = 3^2

Nnesha (nnesha):

we need to get the same base at the numerator and at the denominator.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

2^2=4

Nnesha (nnesha):

yes right \[\huge\rm \frac{ 2 }{ (2^2)^{-3} }\] to solve that you need to know two exponent rules first one is \[\huge\rm (x^m)^n=x^{m · n}\] and then when we divide same bases we should `subtract` their exponents \[\large\rm \frac{ x^b }{x^c }=x^{b-c}\]

OpenStudy (anonymous):

would it be 1/32

Nnesha (nnesha):

nope.

Nnesha (nnesha):

how did you get that ?

OpenStudy (anonymous):

256

Nnesha (nnesha):

hmm how ?..

OpenStudy (anonymous):

is it right

Nnesha (nnesha):

well plz show ur work so i can find where udid a mistake.

OpenStudy (anonymous):

plz tell me if it is right or not first

Nnesha (nnesha):

well that's a guess

OpenStudy (anonymous):

so it is not right is it

OpenStudy (anonymous):

128

Nnesha (nnesha):

\[\huge\rm (2^2)^{-3}\] we should solve it apply the exponent rule \[\huge\rm (x^m )^n =x^{m \times n}\] multiply the exponents

OpenStudy (anonymous):

plz just tell me the answer i only have 2 more minutes to finish it

Nnesha (nnesha):

we can solve this in two minutes i can't give u the answer plz try to understand

OpenStudy (anonymous):

-64

Nnesha (nnesha):

no just multiply the exponents

OpenStudy (anonymous):

5

Nnesha (nnesha):

\[\huge\rm (2^2)^{-3} =2^{2 \times -3}\] base would stay the same

Nnesha (nnesha):

no it's not 5

Nnesha (nnesha):

2 times -3 = ?